OpenStudy (anonymous):

R=log(I/i0) where i0 is the intensity of a certain very small earthquake. The 2010 Chile earthquake had a magnitude of 8.8. What is its intensity (in terms of i0)?

OpenStudy (anonymous):

I will assume R is the magnitude. Then, we have:\[8.8 = \log(\frac{I}{i0})\]Raising both sides to 10 we get:

OpenStudy (anonymous):

\[10^{8.8} = \frac{I}{i_0}\]Multiply both sides by i0 to get:\[10^{8.8}i_0 = I\]

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

No problem :-)